# Modular invariance groups and defect McKay-Thompson series

### v1

@@ 3.2 In $\operatorname{\textsl{SL}}(2,\mathbb{R})$: Atkin-Lehner involutions
In general, since $\operatorname{\textsl{SL}}(2,\mathbb{R})$ is not finitely generated, this is an intractable problem; that said, there is a particular class of $SL(2,\mathbb{R})$ transformations which we are well-motivated to consider: the Atkin-Lehner involutions. Conway and Norton [1] argued that the invariance groups of the McKay-Thompson series have the form $\Gamma_{0}(N)+e,f,g,...$, where $N$ is the level of the invariance group as a congruence subgroup of $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, and $+e,f,g...$ signifies that we may have to adjoin certain cosets of $\Gamma_{0}(N)$, labeled by integers $e,f,g$ which exactly divide $N$ ($e\mathrel{\|}N$ means $e\mid N$ and $(e,N/e)$ are relatively prime). Since we may determine the level of $\Gamma$ even if we only know the invariance group in $SL(2,\mathbb{Z})$, we may assume for the moment that the defect McKay-Thompson series behave similarly, and explicitly check whether any of these cosets fix $Z_{\mathcal{L}}(\tau)$. If not, we know that either the invariance group in $SL(2,\mathbb{R})$ does not extend that in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, or the defect McKay-Thompson series have invariance groups in $SL(2,\mathbb{R})$ of a different form from those of the “normal” McKay-Thompson series.
To perform the checks, we need formulae for the characters appearing in the monstralizer decompositions (2.2). Fortunately, in nearly all of the cases discussed in section 2, one set of characters (those of the “small” RCFT) is given in terms of minimal model characters, and the other set is given either by the action of a vector-valued Hecke operator on the first. Only in the case with monstralizer pair $\quantity(D_{3\text{A}},Fi_{23})$ are the characters related by a modular linear differential equation without a (known) solution in terms of Hecke operators. In all other cases, we may therefore express $Z_{\mathcal{L}}(\tau)$ as an infinite sum. Next, the $\operatorname{\textsl{SL}}(2,\mathbb{R})$-transformed characters are computed from the untransformed ones by Poisson resummation, allowing us to check whether a particular transformation fixes $Z_{\mathcal{L}}(\tau)$. For each series which is not genus zero, we may obtain the level using our method for finding the invariance group in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, and then check each of the associated Atkin-Lehner involutions to see if the series is invariant. Doing so, we do not find that any of the Atkin-Lehner involutions extend the invariance groups in any case (apart from $D_{3A}$, which we cannot check with our methods).
@@ 4 Results
$Z^{W}(q)={}$ $\frac{1}{2}\quantity(\frac{1}{q}+196884q+21493760q^{2}+864299970q^{3}+...)$ (4.2)
 $\mbox{\quad\quad}+\frac{\sqrt{5}}{2}\quantity(\frac{1}{q}-78794q-9307840q^{2}-381508370q^{3}+...).$
@@
$Z^{A}(q)={}$ $\cos(\frac{3\pi}{14})\csc(\frac{\pi}{7})\quantity(\frac{1}{q}-9618984q^{2}+...)$ (4.4)
 $\mbox{\quad\quad}+(36570q+5104206q^{2}+229935232q^{3}+...)$
 $\mbox{\quad\quad}+\frac{1}{2}\csc(\frac{\pi}{14})\quantity(176130q+7900063q^{2}+663713370q^{3}+...)$
 $\mbox{\quad\quad}+\cos(\frac{\pi}{7})\csc(\frac{\pi}{14})\quantity(-88362q-384310822q^{3}+...).$
@@ 5 Conclusions
In the future, we may wish to understand whether the noninvertible lines which are genus zero fit some pattern, as well as how these results fit into the physical interpretation of the genus zero property proposed in [19, 20].
@@ 6 Acknowledgements
This work was supported by the National Science Foundation Grant PHY-2310635. JH thanks S.-H. Shao and G. Moore for helpful conversations.
@@ Appendix A Generators for the invariance groups of high genus
Here we present generators for the invariance groups in table 5 of genus greater than 3, since the information provided in the table is not sufficient to uniquely determine them. The invariance groups of the TDLs of the $\mathbb{Z}_{4}$ parafermion subcategory labeled $X$, $\theta X$, $\theta^{2}X$, and $\theta^{3}X$ are identical, and are generated by
[…]
$\smallmatrixquantity(53&-6\\ 168&-19),\ \smallmatrixquantity(67\hfil&-9\\ 216&-29),\ \smallmatrixquantity(163&-24\\ 360&-53),\smallmatrixquantity(667&-101\\ 984&-149),\ \smallmatrixquantity(25&-4\\ 144&-23),\smallmatrixquantity(265&-49\\ 384&-71),\ \smallmatrixquantity(263&-50\\ 384&-73),$
$\smallmatrixquantity(19&-4\\ 24&-5),\ \smallmatrixquantity(53\hfil&-14\\ 72&-19),\ \smallmatrixquantity(265&-73\\ 432&-119),\ \smallmatrixquantity(139&-39\\ 360&-101),\ \smallmatrixquantity(169&-49\\ 576&-167),\ \smallmatrixquantity(169&-50\\ 240&-71),\ \smallmatrixquantity(53&-19\\ 120&-43),$
$\smallmatrixquantity(73&-27\\ 192&-71),\ \smallmatrixquantity(307\hfil&-120\\ 504&-197),\ \smallmatrixquantity(913&-358\\ 1344&-527),\ \smallmatrixquantity(409&-161\\ 912&-359),\ \smallmatrixquantity(361&-147\\ 528&-215),\ \smallmatrixquantity(317&-130\\ 456&-187),$
$\smallmatrixquantity(121&-50\\ 288&-119),\ \smallmatrixquantity(265\hfil&-121\\ 576&-263),\ \smallmatrixquantity(25&-12\\ 48&-23),\ \smallmatrixquantity(211&-117\\ 312&-173),\ \smallmatrixquantity(77&-43\\ 120&-67),\ \smallmatrixquantity(121&-75\\ 192&-119),\ \smallmatrixquantity(-1&0\\ 0&-1)\big{\}}.$
@@
[…]
$\smallmatrixquantity(281&-13\\ 735&-34),\ \smallmatrixquantity(41\hfil&-2\\ 595&-29),\ \smallmatrixquantity(176&-9\\ 665&-34),\ \smallmatrixquantity(314&-17\\ 665&-36),\ \smallmatrixquantity(71&-4\\ 1225&-69),\ \smallmatrixquantity(314&-19\\ 595&-36),\ \smallmatrixquantity(251&-16\\ 455&-29),$
$\smallmatrixquantity(146&-11\\ 385&-29),\ \smallmatrixquantity(139\hfil&-11\\ 455&-36),\ \smallmatrixquantity(106&-9\\ 1225&-104),\ \smallmatrixquantity(139&-13\\ 385&-36),\ \smallmatrixquantity(426&-41\\ 665&-64),\ \smallmatrixquantity(169&-18\\ 385&-41),\ \smallmatrixquantity(76&-9\\ 245&-29),$
$\smallmatrixquantity(169&-22\\ 315&-41),\ \smallmatrixquantity(244\hfil&-33\\ 525&-71),\ \smallmatrixquantity(29&-5\\ 35&-6),\ \smallmatrixquantity(111&-23\\ 140&-29),\ \smallmatrixquantity(216&-47\\ 455&-99),\ \smallmatrixquantity(244&-55\\ 315&-71),\ \smallmatrixquantity(281&-64\\ 1225&-279),$
$\smallmatrixquantity(64&-15\\ 175&-41),\ \smallmatrixquantity(111\hfil&-29\\ 245&-64),\ \smallmatrixquantity(211&-57\\ 385&-104),\ \smallmatrixquantity(76&-21\\ 105&-29),\ \smallmatrixquantity(99&-29\\ 140&-41),\ \smallmatrixquantity(386&-117\\ 805&-244),\ \smallmatrixquantity(386&-121\\ 1225&-384),$
$\smallmatrixquantity(71&-23\\ 105&-34),\ \smallmatrixquantity(181\hfil&-63\\ 385&-134),\ \smallmatrixquantity(316&-111\\ 595&-209),\ \smallmatrixquantity(181&-64\\ 280&-99),\ \smallmatrixquantity(64&-25\\ 105&-41),\ \smallmatrixquantity(41&-17\\ 70&-29),\ \smallmatrixquantity(99&-43\\ 175&-76),$
$\smallmatrixquantity(561&-256\\ 1225&-559),\ \smallmatrixquantity(246\hfil&-113\\ 455&-209),\ \smallmatrixquantity(274&-131\\ 525&-251),\ \smallmatrixquantity(414&-199\\ 595&-286),\ \smallmatrixquantity(356&-187\\ 455&-239),\ \smallmatrixquantity(134&-85\\ 175&-111),\ \smallmatrixquantity(-1&0\\ 0&-1)\big{\}}.$

### v2

@@ 3.2 In $\operatorname{\textsl{SL}}(2,\mathbb{R})$: Atkin-Lehner involutions
In general, since $\operatorname{\textsl{SL}}(2,\mathbb{R})$ is not finitely generated, this is an intractable problem; that said, there is a particular class of $SL(2,\mathbb{R})$ transformations which we are well-motivated to consider: the Atkin-Lehner involutions. Conway and Norton [1] argued that the invariance groups of the McKay-Thompson series have the form $\Gamma_{0}(N)+e,f,g,$..., where $N$ is the level of the invariance group as a congruence subgroup of $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, and +e,f,g+e,f,g... signifies that we may have to adjoin certain cosets of $\Gamma_{0}(N)$, labeled by integers $e,f,g$ which exactly divide $N$ ($e\mathrel{\|}N$ means $e\mid N$ and $(e,N/e)$ are relatively prime). Since we may determine the level of $\Gamma$ even if we only know the invariance group in $SL(2,\mathbb{Z})$, we may assume for the moment that the defect McKay-Thompson series behave similarly, and explicitly check whether any of these cosets fix $Z_{\mathcal{L}}(\tau)$. If not, we know that either the invariance group in $SL(2,\mathbb{R})$ does not extend that in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, or the defect McKay-Thompson series have invariance groups in $SL(2,\mathbb{R})$ of a different form from those of the “normal” McKay-Thompson series.
To perform the checks, we need formulae for the characters appearing in the monstralizer decompositions (2.1). Fortunately, in nearly all of the cases discussed in section 2, one set of characters (those of the “small” RCFT) is given in terms of minimal model characters, and the other set is given by the action of a vector-valued Hecke operator on the first. Only in the case with monstralizer pair $\quantity(D_{3\text{A}},Fi_{23})$ are the characters related by a modular linear differential equation without a (known) solution in terms of Hecke operators. In all other cases, we may therefore express $Z_{\mathcal{L}}(\tau)$ as an infinite sum. Next, the $\operatorname{\textsl{SL}}(2,\mathbb{R})$-transformed characters are computed from the untransformed ones by Poisson resummation, allowing us to check whether a particular transformation fixes $Z_{\mathcal{L}}(\tau)$. For each series which is not genus zero, we may obtain the level using our method for finding the invariance group in $\operatorname{\textsl{SL}}(2,\mathbb{Z})$, and then check each of the associated Atkin-Lehner involutions to see if the series is invariant. Doing so, we do not find that any of the Atkin-Lehner involutions extend the invariance groups in any case (apart from $D_{3A}$, which we cannot check with our methods).
@@ 4 Results
[…]
@@
[…]
@@ 5 Conclusions
In the future, we may wish to understand whether the non-invertible lines which are genus zero fit some pattern, as well as how these results fit into the physical interpretation of the genus zero property proposed in [21, 22].
@@ 6 Acknowledgements
This work was supported by the National Science Foundation Grant PHY-2310635. JH thanks S.-H. Shao and G. Moore for helpful conversations. We thank one of the referees for several helpful remarks.
@@ Appendix A Generators for the invariance groups of high genus
Here we present generators for the invariance groups in table 5 of genus greater than 3, since the information provided in the table is not sufficient to uniquely determine them.
The invariance group of the TDL of the $\mathbb{Z}_{3}$ parafermion subcategory labeled $WN$ is generated by
$\big{\{}{}$ $\quantity(\smallmatrixquantity{1&1\\ 0&1}),\quantity(\smallmatrixquantity{109&-5\\ 240&-11}),\quantity(\smallmatrixquantity{61&-3\\ 1200&-59}),\quantity(\smallmatrixquantity{641&-33\\ 2700&-139}),\quantity(\smallmatrixquantity{1019&-53\\ 3480&-181}),\quantity(\smallmatrixquantity{391&-21\\ 540&-29}),\quantity(\smallmatrixquantity{209&-12\\ 540&-31}),\quantity(\smallmatrixquantity{989&-60\\ 1500&-91}),$ (A.1)
$\quantity(\smallmatrixquantity{221&-14\\ 300&-19}),\quantity(\smallmatrixquantity{191&-13\\ 720&-49}),\quantity(\smallmatrixquantity{541&-38\\ 840&-59}),\quantity(\smallmatrixquantity{149&-11\\ 420&-31}),\quantity(\smallmatrixquantity{1261&-99\\ 2280&-179}),\quantity(\smallmatrixquantity{1499&-119\\ 2280&-181}),\quantity(\smallmatrixquantity{61&-5\\ 720&-59}),$
$\quantity(\smallmatrixquantity{439&-38\\ 1860&-161}),\quantity(\smallmatrixquantity{581&-51\\ 900&-79}),\quantity(\smallmatrixquantity{109&-10\\ 120&-11}),\quantity(\smallmatrixquantity{161&-17\\ 180&-19}),\quantity(\smallmatrixquantity{631&-72\\ 780&-89}),\quantity(\smallmatrixquantity{1379&-160\\ 1560&-181}),\quantity(\smallmatrixquantity{421&-49\\ 3600&-419}),$
$\quantity(\smallmatrixquantity{401&-48\\ 660&-79}),\quantity(\smallmatrixquantity{299&-38\\ 480&-61}),\quantity(\smallmatrixquantity{139&-18\\ 780&-101}),\quantity(\smallmatrixquantity{581&-76\\ 1980&-259}),\quantity(\smallmatrixquantity{659&-87\\ 2280&-301}),\quantity(\smallmatrixquantity{139&-19\\ 300&-41}),\quantity(\smallmatrixquantity{169&-25\\ 480&-71}),$
$\quantity(\smallmatrixquantity{251&-38\\ 720&-109}),\quantity(\smallmatrixquantity{31&-5\\ 180&-29}),\quantity(\smallmatrixquantity{139&-26\\ 540&-101}),\quantity(\smallmatrixquantity{101&-19\\ 420&-79}),\quantity(\smallmatrixquantity{691&-132\\ 780&-149}),\quantity(\smallmatrixquantity{539&-104\\ 1560&-301}),\quantity(\smallmatrixquantity{619&-120\\ 1140&-221}),$
$\quantity(\smallmatrixquantity{289&-57\\ 360&-71}),\quantity(\smallmatrixquantity{191&-39\\ 240&-49}),\quantity(\smallmatrixquantity{349&-73\\ 1200&-251}),\quantity(\smallmatrixquantity{151&-32\\ 420&-89}),\quantity(\smallmatrixquantity{659&-142\\ 840&-181}),\quantity(\smallmatrixquantity{781&-169\\ 3600&-779}),\quantity(\smallmatrixquantity{229&-50\\ 600&-131}),$
$\quantity(\smallmatrixquantity{79&-18\\ 180&-41}),\quantity(\smallmatrixquantity{211&-49\\ 900&-209}),\quantity(\smallmatrixquantity{1171&-277\\ 3420&-809}),\quantity(\smallmatrixquantity{409&-106\\ 1200&-311}),\quantity(\smallmatrixquantity{991&-257\\ 1500&-389}),\quantity(\smallmatrixquantity{589&-154\\ 960&-251}),\quantity(\smallmatrixquantity{571&-153\\ 780&-209}),$
$\quantity(\smallmatrixquantity{329&-89\\ 780&-211}),\quantity(\smallmatrixquantity{559&-154\\ 1020&-281}),\quantity(\smallmatrixquantity{661&-183\\ 1080&-299}),\quantity(\smallmatrixquantity{329&-92\\ 540&-151}),\quantity(\smallmatrixquantity{991&-286\\ 1140&-329}),\quantity(\smallmatrixquantity{1259&-365\\ 2280&-661}),\quantity(\smallmatrixquantity{1139&-333\\ 1440&-421}),$
$\quantity(\smallmatrixquantity{169&-50\\ 240&-71}),\quantity(\smallmatrixquantity{41&-13\\ 60&-19}),\quantity(\smallmatrixquantity{4921&-1681\\ 14400&-4919}),\quantity(\smallmatrixquantity{3001&-1026\\ 4560&-1559}),\quantity(\smallmatrixquantity{449&-154\\ 1140&-391}),\quantity(\smallmatrixquantity{311&-107\\ 840&-289}),\quantity(\smallmatrixquantity{449&-156\\ 780&-271}),$
$\quantity(\smallmatrixquantity{331&-121\\ 900&-329}),\quantity(\smallmatrixquantity{419&-158\\ 480&-181}),\quantity(\smallmatrixquantity{1009&-383\\ 2400&-911}),\quantity(\smallmatrixquantity{1301&-495\\ 2100&-799}),\quantity(\smallmatrixquantity{389&-152\\ 540&-211}),\quantity(\smallmatrixquantity{581&-228\\ 660&-259}),\quantity(\smallmatrixquantity{71&-29\\ 120&-49}),$
$\quantity(\smallmatrixquantity{349&-146\\ 600&-251}),\quantity(\smallmatrixquantity{1861&-781\\ 3000&-1259}),\quantity(\smallmatrixquantity{509&-216\\ 780&-331}),\quantity(\smallmatrixquantity{919&-420\\ 1140&-521}),\quantity(\smallmatrixquantity{649&-298\\ 1200&-551}),\quantity(\smallmatrixquantity{2971&-1366\\ 4500&-2069}),\quantity(\smallmatrixquantity{1979&-911\\ 3000&-1381}),$
$\quantity(\smallmatrixquantity{31&-15\\ 60&-29}),\quantity(\smallmatrixquantity{259&-139\\ 300&-161}),\quantity(\smallmatrixquantity{661&-363\\ 1200&-659}),\quantity(\smallmatrixquantity{869&-481\\ 1140&-631}),\quantity(\smallmatrixquantity{139&-78\\ 180&-101}),\quantity(\smallmatrixquantity{569&-329\\ 780&-451}),\quantity(\smallmatrixquantity{749&-454\\ 1140&-691}),$
$\quantity(\smallmatrixquantity{469&-290\\ 600&-371}),\quantity(\smallmatrixquantity{551&-347\\ 840&-529}),\quantity(\smallmatrixquantity{571&-361\\ 900&-569}),\quantity(\smallmatrixquantity{331&-212\\ 420&-269}),\quantity(\smallmatrixquantity{409&-265\\ 480&-311}),\quantity(\smallmatrixquantity{611&-398\\ 720&-469}),\quantity(\smallmatrixquantity{1259&-824\\ 1560&-1021}),$
$\quantity(\smallmatrixquantity{929&-612\\ 1140&-751}),\quantity(\smallmatrixquantity{341&-259\\ 420&-319}),\quantity(\smallmatrixquantity{691&-529\\ 900&-689}),\quantity(\smallmatrixquantity{679&-558\\ 780&-641}),\quantity(\smallmatrixquantity{151&-125\\ 180&-149}),\quantity(\smallmatrixquantity{-1&0\\ 0&-1})\big{\}}.$
The invariance groups of the TDLs of the $\mathbb{Z}_{4}$ parafermion subcategory labeled $X$, $\theta X$, $\theta^{2}X$, and $\theta^{3}X$ are identical, and are generated by
$\big{\{}$ $\quantity(\smallmatrixquantity{1&1\\ 0&1}),\ \quantity(\smallmatrixquantity{43&-3\\ 72&-5}),\ \quantity(\smallmatrixquantity{25&-2\\ 288&-23}),\ \quantity(\smallmatrixquantity{73&-7\\ 240&-23}),\ \quantity(\smallmatrixquantity{605&-62\\ 888&-91}),\quantity(\smallmatrixquantity{241&-25\\ 2304&-239}),\quantity(\smallmatrixquantity{265&-28\\ 672&-71}),$ (A.2)
$\quantity(\smallmatrixquantity{53&-6\\ 168&-19}),\ \quantity(\smallmatrixquantity{67&-9\\ 216&-29}),\ \quantity(\smallmatrixquantity{163&-24\\ 360&-53}),\quantity(\smallmatrixquantity{667&-101\\ 984&-149}),\ \quantity(\smallmatrixquantity{25&-4\\ 144&-23}),\quantity(\smallmatrixquantity{265&-49\\ 384&-71}),\ \quantity(\smallmatrixquantity{263&-50\\ 384&-73}),$
$\quantity(\smallmatrixquantity{19&-4\\ 24&-5}),\ \quantity(\smallmatrixquantity{53&-14\\ 72&-19}),\ \quantity(\smallmatrixquantity{265&-73\\ 432&-119}),\ \quantity(\smallmatrixquantity{139&-39\\ 360&-101}),\ \quantity(\smallmatrixquantity{169&-49\\ 576&-167}),\ \quantity(\smallmatrixquantity{169&-50\\ 240&-71}),\ \quantity(\smallmatrixquantity{53&-19\\ 120&-43}),$
$\quantity(\smallmatrixquantity{73&-27\\ 192&-71}),\ \quantity(\smallmatrixquantity{307&-120\\ 504&-197}),\ \quantity(\smallmatrixquantity{913&-358\\ 1344&-527}),\ \quantity(\smallmatrixquantity{409&-161\\ 912&-359}),\ \quantity(\smallmatrixquantity{361&-147\\ 528&-215}),\ \quantity(\smallmatrixquantity{317&-130\\ 456&-187}),$
$\quantity(\smallmatrixquantity{121&-50\\ 288&-119}),\ \quantity(\smallmatrixquantity{265&-121\\ 576&-263}),\ \quantity(\smallmatrixquantity{25&-12\\ 48&-23}),\ \quantity(\smallmatrixquantity{211&-117\\ 312&-173}),\ \quantity(\smallmatrixquantity{77&-43\\ 120&-67}),\ \quantity(\smallmatrixquantity{121&-75\\ 192&-119}),\ \quantity(\smallmatrixquantity{-1&0\\ 0&-1})\big{\}}.$
@@
$\big{\{}{}$ $\quantity(\smallmatrixquantity{1&1\\ 0&1}),\ \quantity(\smallmatrixquantity{29&-1\\ 175&-6}),\ \quantity(\smallmatrixquantity{111&-4\\ 805&-29}),\ \quantity(\smallmatrixquantity{531&-20\\ 770&-29}),\ \quantity(\smallmatrixquantity{76&-3\\ 735&-29}),\ \quantity(\smallmatrixquantity{559&-23\\ 875&-36}),\ \quantity(\smallmatrixquantity{71&-3\\ 805&-34}),\ \quantity(\smallmatrixquantity{321&-14\\ 665&-29}),$ (A.3)
$\quantity(\smallmatrixquantity{281&-13\\ 735&-34}),\ \quantity(\smallmatrixquantity{41&-2\\ 595&-29}),\ \quantity(\smallmatrixquantity{176&-9\\ 665&-34}),\ \quantity(\smallmatrixquantity{314&-17\\ 665&-36}),\ \quantity(\smallmatrixquantity{71&-4\\ 1225&-69}),\ \quantity(\smallmatrixquantity{314&-19\\ 595&-36}),\ \quantity(\smallmatrixquantity{251&-16\\ 455&-29}),$
$\quantity(\smallmatrixquantity{146&-11\\ 385&-29}),\ \quantity(\smallmatrixquantity{139&-11\\ 455&-36}),\ \quantity(\smallmatrixquantity{106&-9\\ 1225&-104}),\ \quantity(\smallmatrixquantity{139&-13\\ 385&-36}),\ \quantity(\smallmatrixquantity{426&-41\\ 665&-64}),\ \quantity(\smallmatrixquantity{169&-18\\ 385&-41}),\ \quantity(\smallmatrixquantity{76&-9\\ 245&-29}),$
$\quantity(\smallmatrixquantity{169&-22\\ 315&-41}),\ \quantity(\smallmatrixquantity{244&-33\\ 525&-71}),\ \quantity(\smallmatrixquantity{29&-5\\ 35&-6}),\ \quantity(\smallmatrixquantity{111&-23\\ 140&-29}),\ \quantity(\smallmatrixquantity{216&-47\\ 455&-99}),\ \quantity(\smallmatrixquantity{244&-55\\ 315&-71}),\ \quantity(\smallmatrixquantity{281&-64\\ 1225&-279}),$
$\quantity(\smallmatrixquantity{64&-15\\ 175&-41}),\ \quantity(\smallmatrixquantity{111&-29\\ 245&-64}),\ \quantity(\smallmatrixquantity{211&-57\\ 385&-104}),\ \quantity(\smallmatrixquantity{76&-21\\ 105&-29}),\ \quantity(\smallmatrixquantity{99&-29\\ 140&-41}),\ \quantity(\smallmatrixquantity{386&-117\\ 805&-244}),\ \quantity(\smallmatrixquantity{386&-121\\ 1225&-384}),$
$\quantity(\smallmatrixquantity{71&-23\\ 105&-34}),\ \quantity(\smallmatrixquantity{181&-63\\ 385&-134}),\ \quantity(\smallmatrixquantity{316&-111\\ 595&-209}),\ \quantity(\smallmatrixquantity{181&-64\\ 280&-99}),\ \quantity(\smallmatrixquantity{64&-25\\ 105&-41}),\ \quantity(\smallmatrixquantity{41&-17\\ 70&-29}),\ \quantity(\smallmatrixquantity{99&-43\\ 175&-76}),$
$\quantity(\smallmatrixquantity{561&-256\\ 1225&-559}),\ \quantity(\smallmatrixquantity{246&-113\\ 455&-209}),\ \quantity(\smallmatrixquantity{274&-131\\ 525&-251}),\ \quantity(\smallmatrixquantity{414&-199\\ 595&-286}),\ \quantity(\smallmatrixquantity{356&-187\\ 455&-239}),\ \quantity(\smallmatrixquantity{134&-85\\ 175&-111}),\ \quantity(\smallmatrixquantity{-1&0\\ 0&-1})\big{\}}.$